OpenStudy (anonymous):

lim x→infinity of [ (sqrt of 4x^6 − x) / (x^3 + 9)]

OpenStudy (anonymous):

\[\lim_{x \rightarrow \infty} \frac{ \sqrt{4x^6-x} }{ x^3+9 }\]

OpenStudy (anonymous):

from your eyeballs \[\frac{\sqrt{4x^6}}{x^3}=\frac{2x^3}{x^3}=2\]

OpenStudy (anonymous):

Oh so the whole limit is \[\lim_{x \rightarrow \infty} of 2\]

OpenStudy (anonymous):

is the answe 2?

OpenStudy (anonymous):

yes, the answer is 2

OpenStudy (anonymous):

oh wow I didnt even see that it was equivalent to 2! thanks

OpenStudy (anonymous):

yw, i did cheat a bit by ignoring the unimportant terms at the end